A sofa is 6 feet long. In a scale drawing, the sofa is 3 inches long. Find the scale of the drawing without units.

Found 2 solutions by stanbon, drj:
Answer by stanbon(75887)

A sofa is 6 feet long. In a scale drawing, the sofa is 3 inches long. Find the scale of the drawing without units.
-----
6 feet = 72 inches
---
72/3 = 24
Ratio: 24:1
======================
Cheers,
Stan H.

Answer by drj(1380)

A sofa is 6 feet long. In a scale drawing, the sofa is 3 inches long. Find the scale of the drawing without units.

Given the sofa is 6 feet long, the scale drawing 3 inches long. Let s be the scale ratio without units. Note 1 foot = 12 inches then 1/4 feet is 3 inches.

I'll give two scale factors depending how you want the answer.

APPROACH 1. Hence, s=6/3 or s=1/2. This means at if the length is 3 feet then the scale drawing is 3*(1/2)=1.5 inches.

So this is a scale conversion of feet/inches (so this is not unitless)

APPROACH 2. We can change the scale factor by changing the denominator to 1/4=3 inches. So

s=6/(1/4)=24. So the scaling is 24 to 1. So 6 feet/24= 1/4 feet or 3 inches in the scale drawing. So this is unitless scale s since the conversion is feet/feet.

I hope the above steps were helpful.

For free Step-By-Step Videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra or for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J